OpenStudy (anonymous):

Which shows solutions to x + 2y = 8 if x and y must be whole numbers. A. {( 2,4), (1,6), (0,8)} B. {(8,0), (6,1), (4,2)} C. {(0,8), (1,5), (2,4)} D. {(0,8), (1,7), (2,4)}

DivineSolar (divinesolar):

I would say A, but wait for others answers, and arent they already whole numbers X3

DivineSolar (divinesolar):

Heres the graph for it if you would like to look it over.

OpenStudy (anonymous):

http://openstudy.com/study#/updates/528e6895e4b0424725715035 ok but i found this i dont know if its accurate

OpenStudy (cwrw238):

plug in the 3 sets of values and see which ones fit its not A

DivineSolar (divinesolar):

Okay.

OpenStudy (cwrw238):

(2,4) in A gives 2 + 4(2) = 2+8 = 10

DivineSolar (divinesolar):

Wouldnt that be C then?

OpenStudy (anonymous):

or B?

OpenStudy (cwrw238):

Not C

OpenStudy (cwrw238):

try B 8 + 2(0) = 8 - the first one fits now try (6,1) and (4,2)

OpenStudy (cwrw238):

(6,1) means x=6 and y=1

DivineSolar (divinesolar):

Oh that makes so much more sense lol.

OpenStudy (anonymous):

so its B

OpenStudy (cwrw238):

so 6 + 2(1) = /

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

thank you just to make sure are you 100% sure because i really need this right

OpenStudy (anonymous):

@cwrw238

OpenStudy (cwrw238):

i'm sure

OpenStudy (anonymous):

thank you